Employees of a publishing company received an increase in salary of 8% plus a bonus of $800 Let S(x) represent the new salary in terms of the previous salary x. Find the value of S (24,000)

Found 2 solutions by Cromlix, robertb:
Answer by Cromlix(4381)

Hi there,
S(x) = 1.08x + 800
S(24,000)= 1.08(24,000) + 800
S(24,000) = $26,720
Hope this helps :-)

Answer by robertb(5830)

S(x) = 1.08x+800, where x = previous salary.
==> S(24,000) = 1.08*24,000+800 = 26,720.